Bob plays a game in which he rolls a pair of fair 6-sided dice once. He adds the number of dots on both dice to create a sum. He

  • loses $1 if he rolls a sum less than 5

  • doesn’t win or lose anything if he rolls a sum that is higher than 4, but less than 8

  • wins $1 if he rolls a sum that’s higher than 7 but less than 12

  • wins $2 if he rolls a sum of 12

a. If X is a random variable indicating Bob’s earning, write the PMF of X in a table.

b. If Y = X2 + 1, write a table for the PMF of Y.
Answer #1

let A is outcome on dice ,

a)

below is pmf of X:

P(X=-1)=P(getting a sum less then 5)=P(A=2)+P(A=3)+P(A=4)=(1/36)+(2/36)+(3/36)=1/6

P(X=0)=P(5<=A<=7)=P(A=5)+P(A=6)+P(A=7)=(4/36)+(5/36)+(6/36)=15/36=5/12

P(X=1)=P(A=8)+P(A=9)+P(A=10)+P(A=11)=(5/36)+(4/36)+(3/36)+(2/36)=14/36=7/18

P(X=2)=P(A=12)=1/36

b)

Y = X^2 +1

P(Y=1)=P(X=0)=5/12

P(Y=2)=P(X=-1)+P(X=1)=1/6+7/18=20/36=5/9

P(Y=5)=P(X=2)=1/36